# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v2

@@ 4.5 Distribution Alignment Normalization
A notable challenge during finetuning is the misalignment of intermediate sample’s marginal distribution $q(\mathbf{z}_{t})$ between DPBridge and diffusion backbone, due to the difference in the forward transition kernel. This misalignment will weaken the model’s inherent denoising ability, thus limiting the effectiveness of knowledge transfer. To address this issue, we propose a distribution alignment normalization technique to fill in the gap between the two distributions. Specifically, we normalize the intermediate latents computed from Equation 3 to match the form of a standard diffusion process:
@@
Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned. Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone. An overview of DPBridge’s training pipeline shown in Algorithm 1.
4.6 Accelerate Inference
Similar to the idea of DDIM (Song et al., 2020a), the inference process of DPBridge can be accelerated by utilizing a non-Markovian process while maintaining the same marginal distributions for all intermediate samples along the sampling path. Specifically, we can sample on a subset of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$ and the number of sampling steps $N$ can be set independently from the original timesteps $T$ on which the bridge model is trained. The accelerated inference process can be thus described as:
where $g_{n}$ is the variance parameter that controls the stochasticity behavior of the reverse sampling process. We note that when $g_{n}=\tilde{\sigma}_{t}^{2}$ which is calculated from Equation 5, the process becomes a Markovian bridge. The basic inference process of DPBridge is summarized in Algorithm 2. If adopt acceleration, one can utilize Equation 13 during implementation.
@@ 5 Experiments
We conduct evaluation mainly on two representative dense prediction tasks, i.e., depth estimation and surface normal prediction. We’ll start by a brief introduction of experiment settings as well as implementation details, followed by quantitative and qualitative analysis. Lastly, we conduct detailed ablation studies and investigate the influence each design component respectively.
@@ 5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim (Roberts et al., 2021) and Virtual KITTI (Cabon et al., 2020), which cover both indoor and outdoor scenes, for training. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 (Silberman et al., 2012), KITTI (Geiger et al., 2012), ScanNet (Dai et al., 2017), ETH3D (Schops et al., 2017) and DIODE (Vasiljevic et al., 2019). For surface normal prediction, we additionally incorporate iBims (Koch et al., 2018) and Sintel (Butler et al., 2012) datasets. For fair comparison, we use the same validation split as in previous methods (Ke et al., 2024).
@@
5.2 Implementation Details
We implement DPBridge based on Stable Diffusion v1.5 (SD1.5) and v2.1-base (SD2.1) (Rombach et al., 2022) and report their results respectively to shown compatibility with different backbones. During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
@@ 5.2 Implementation Details
DiverseDepth (Yin et al., 2020) 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.4
MiDaS (Ranftl et al., 2020) 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
LeReS (Yin et al., 2021) 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
Omnidata (Eftekhar et al., 2021) 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
HDN (Zhang et al., 2022) 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
DPT (Ranftl et al., 2021) 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 6.9
Depth Anything V2 (Yang et al., 2024) 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
Marigold (Ke et al., 2024) 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 5.8
DepthFM (Gui et al., 2024) 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
Geowizard (Fu et al., 2025) 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.1
E2E-FT (Garcia et al., 2024) 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.4
I2SB (Liu et al., 2023) 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1 13.9
BBDM (Li et al., 2023) 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7 14.1
DPBridge (SD1.5) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6 6.1
DPBridge (SD2.1) 4.4 97.6 9.8 91.6 6.9 94.9 4.5 97.4 23.8 84.0 2.6
Table 1: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@
EESNU (Bae et al., 2021) 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.5
Omnidata v1 (Eftekhar et al., 2021) 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.1
Omnidata v2 (Kar et al., 2022) 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 5.6
DSINE (Bae & Davison, 2024) 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.2
Marigold (Ke et al., 2024) 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 5.8
Geowizard (Fu et al., 2025) 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.1
E2E-FT (Garcia et al., 2024) 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
DPBridge (SD1.5) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4 7.3
@@
Table 2: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@ 5.3 Quantitative and Qualitative Results
Table 1 and Table 2 summarize the quantitative metrics of DPBridge compared with state-of-the-art (SOTA) methods from both discriminative and generative fields, all in zero-shot manner. In depth estimation, DPBridge consistently matches SOTA results, while demonstrating robust generalization across datasets. It achieves 4.4 AbsRel, 97.2 % $\delta_{1}$ on NYUv2, and 4.5 AbsRel, 97.4% $\delta_{1}$ on ScanNet. Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 14.8 and $11.25^{\circ}$ accuracy of $65.3\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 17.1, $11.25^{\circ}:68.0\%$) and Sintel (mean: 25.0, $11.25^{\circ}:46.3\%$). These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks. Notably, when compared with bridge models like I2SB and BBDM that are trained from scratch, we exceed their performance by a large margin, justifying our proposal that adapting prior knowledge to bridge construction indeed benefits the perception performance. In terms of diffusion-based methods, DPBridge can also achieve the best result, showcasing that adopting bridge process to handle dense prediction tasks is indeed advantageous over standard diffusion framework.
We further provide qualitative examples for both tasks in comparison with SOTA diffusion-based methods (Ke et al., 2024; Fu et al., 2025; Garcia et al., 2024), which are most closely related to our work. Figure 3 shows that E2E-FT tends to produce blurry and smooth predictions, while Geowizard fails to provide accurate estimation in relatively long distance range. In contrast, DPBridge perform well in both indoor and outdoor scenes, capturing fine-level details in both near and faraway regions, like the contour of leaves, fences, cars, computers or windows. Figure 4 presents visualizations on in-the-wild images for surface normal prediction task, where we can see that DPBridge is able to predict high-quality surface normals equipped with fine-grained textures.
Figure 3: Qualitative comparison of diffusion-based depth estimation methods on KITTI and ETH3D dataset. DPBridge is able to generalize in both indoor and outdoor scenes, successively capturing fine-grained details such as leaves and structures on the fence (KITTI).
Figure 4: Qualitative comparison between DPBridge (above) and Marigold (below) on normal estimation results, DPBridge is able to capture more textures from the photos in-the-wild.
@@
Figure 5: Ablation of sampling steps. The performance improves as the number of sampling steps increases, while we observe saturation after 10 steps.
@@ 5.4 Ablation Studies
Finetuning VS Training From Scratch. To assess whether adapting pretrained diffusion backbones improves perception performance, we conduct comparison between finetuning and training from scratch paradigm. As shown in the first row of Table 3, fine-tuning significantly outperforms training from scratch under identical experimental settings. This difference is due to two factors: first, foundation models trained on large, diverse datasets. They capture universal visual priors, aiding knowledge transfer to dense prediction tasks. Second, training from scratch requires more iterations for convergence. In addition, from the last two rows of Table 1 and Table 2, we can see that finetuning on SD2.1 exceeds that on SD1.5, suggesting that the quality of prior knowledge will also influence the performance. These observations indicate that building DPBridge on superior pretrained backbones is indeed advantageous.
Distribution Alignment Normalization. Next, we evaluate whether aligning the marginal distributions between diffusion bridge process and standard diffusion process is beneficial. As shown in the second and fourth rows of Table 3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
Image Consistency Loss. Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Table 3 shows that while effective, its impact is less significant than that of distribution alignment normalization. This is probably because it primarily influences regions where masked and valid areas overlap, which constitute only a relatively small proportion of the data. However, when integrated with other design components, the overall performance is significantly improved.
Number of Sampling Steps. We evaluate performance with varying numbers of sampling steps during inference. As shown in Figure 5, increasing sampling steps initially improves perception performance rapidly and reach saturation after 10 steps. Notably, compared to the results in Table 1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
@@ 6 Conclusion
In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks through generative pipeline. We establish direct mapping between input images and corresponding signal maps in the latent space based on fully-tractable diffusion bridge process. In this way, the prior distribution is no longer constrained to uninformative noise, but instead can be chosen as more structurally correlated ones. In addition, we introduce effective finetuning strategies to adapt large pretrained diffusion backbone to our bridge model construction, leveraging its rich prior knowledge to boost training speed and achieve remarkable perception performance. Future research directions include deriving other parameterization of diffusion bridge process, and extending to more dense prediction tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and we wish to spark further research on image generation and perception. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4.1. Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation 5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$. The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Equation 3:
@@
which is the same as in Equation 8 and Equation 9. This concludes our derivation.
Appendix B Experiments
B.1 Network Architecture
The basic structure of our network follows the original design of Stable Diffusion. To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate samples and input image along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by (Ke et al., 2024), we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
B.2 Additional Quantitative Results

### v3

@@ 4.5 Distribution Alignment Normalization
A key challenge during finetuning is the misalignment between the marginal distribution $q(\mathbf{z}_{t})$ of intermediate samples in DPBridge and that of the diffusion backbone, due to the difference in the forward transition kernels. This misalignment weakens the model’s denoising ability and hinders effective knowledge transfer. To address this issue, we propose a distribution alignment normalization technique that transforms intermediate latents computed from our bridge process Equation 3 to match the distributional form of a standard diffusion process.
@@
Since the square sum of coefficients is equal to 1, the normalized bridge process $\{\mathbf{z}_{t}^{\prime}:t\in[0,T]\}$ is the same as a standard VP process, making the marginal distribution of these two aligned. Therefore by feeding the normalized sample $\mathbf{z}_{t}^{\prime}$ in the network $\boldsymbol{\epsilon}_{\theta}$, we can fully exploit the denoising capability from pretrained diffusion backbone. We note that a similar strategy was independently proposed in FrameBridge [52], where it is applied in the context of image-to-video generation. Our method, developed concurrently, is specifically designed to facilitate adaptation in dense prediction tasks under the DPBridge framework. An overview of the full training and inference procedures are summarized in Algorithm 1 Algorithm 2. For more derivation details, please refer to Appendix A.
@@ 5 Experiments
We conduct evaluation primarily on two representative dense prediction tasks: depth estimation and surface normal prediction. We begin with a brief overview of experimental settings, followed by both quantitative and qualitative analyses. Next, we present ablation studies to investigate the contribution of each design component. Implementation details are presented in Appendix D for reference.
@@ 5.1 Experiments Setup
Datasets. The basic setting for depth estimation and surface normal prediction experiments are similar. We use two synthetic datasets, HyperSim [40] and Virtual KITTI [6], for training, which cover both indoor and outdoor scenes with high-quality ground-truth annotations. To comprehensively evaluate the model performance and demonstrate its generalization capability, we conduct zero-shot evaluations on real-world datasets. For depth estimation, this includes NYUv2 [44], KITTI [18], ScanNet [13], ETH3D [43] and DIODE [50]. For surface normal prediction, we additionally incorporate iBims [29] and Sintel [5] datasets. For fair comparison, we use the same validation split as in previous methods [28].
@@
DiverseDepth [58] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1 12.2
MiDaS [38] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5 11.8
LeReS [59] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6 9.5
Omnidata [15] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2 9.2
HDN [61] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0 7.2
DPT [39] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8 7.0
Depth Anything V2 [56] 4.4 97.9 7.5 94.8 13.2 84.2 4.4 98.0 6.5 95.4 2.3
Marigold [28] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2 6.0
DepthFM [19] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0 3.8
Geowizard [16] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1 6.2
E2E-FT [17] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6 3.6
I2SB [32] 23.5 63.6 35.1 46.4 53.4 40.5 21.1 67.4 55.5 58.2 14.0
BBDM [31] 33.2 49.6 58.6 25.6 52.0 40.6 25.2 61.2 70.1 46.9 15.2
DDBM [65] 26.4 62.2 55.1 27.6 69.5 33.8 23.2 65.6 52.3 56.9 14.5
DPBridge (SD1.5) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6 5.4
DPBridge (SD2.1) 4.4 97.6 9.8 91.6 6.9 94.9 4.5 97.4 23.8 84.0 2.7
Table 1: Quantitative comparison on affine-invariant depth estimation benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
@@
EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5 6.8
Omnidata v1 [15] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4 8.8
Omnidata v2 [27] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7 6.2
DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5 3.5
Marigold [28] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9 6.4
Geowizard [16] 17.0 56.5 15.4 61.6 13.0 65.3 40.4 13.2 4.2
E2E-FT [17] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3 2.1
StableNormal [57] 17.7 55.9 15.8 61.2 17.2 66.6 - - 5.0
DPBridge (SD1.5) 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4 8.0
@@
Table 2: Quantitative comparison on surface normal prediction benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The best and second-best results are highlighted in bold and underline formats.
Input Image
DPBridge
Marigold
Geowizard
DepthFM
Figure 4: Qualitative comparison with diffusion-based depth estimation methods on in-the-wild images. DPBridge can generate depth maps with sharper boundaries and more coherent scene geometry.
Input Image
DPBridge
Marigold
Geowizard
StableNormal
Figure 5: Qualitative comparison with diffusion-based surface normal prediction methods on in-the-wild images. DPBridge’s results exhibit more fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat.
@@ 5.2 Quantitative and Qualitative Results
Table 1 and Table 2 summarize the quantitative results of DPBridge on zero-shot benchmarks across depth estimation and surface normal prediction tasks. We evaluate two variants of our model—DPBridge (SD1.5) and DPBridge (SD2.1)—built on top of Stable Diffusion v1.5 and v2.1-base [41], respectively. For depth estimation, DPBridge (SD2.1) achieves an AbsRel of 4.4 and $\delta_{1}$ accuracy of 97.6% on NYUv2, and 4.5 AbsRel / 97.4% $\delta_{1}$ on ScanNet, both among the best in class. Across all five datasets, it achieves an average ranking of 2.7—second only to Depth Anything V2 (rank 2.3), which is trained on over 1,000 times more data. For surface normal prediction, DPBridge also delivers state-of-the-art performance with a mean angular error of 14.8° and $65.3\%$ accuracy under $11.25^{\circ}$ on NYUv2, outperforming both feed-forward and diffusion-based methods. It also maintains strong results on iBims-1 ($17.1$, $68.0\%$) and Sintel ($25.0$, $46.3\%$). These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks. Notably, compared to prior bridge models such as I2SB, BBDM, and DDBM, DPBridge reduces average error by a large margin (e.g., 4.4 vs. 23.5 AbsRel in NYUv2 depth), underscoring the benefits and importance of incorporating pretrained visual priors to bridge construction. Furthermore, among diffusion-based approaches, DPBridge consistently achieves the best or second-best scores, validating our core hypothesis that bridge formulation provides clear advantages over standard diffusion frameworks in dense prediction tasks, and combining bridge modeling with pretrained visual prior offers a more efficient and generalizable framework.
Figure 5 and Figure 5 present qualitative comparisons of DPBridge with SOTA diffusion-based methods on in-the-wild images for depth estimation and surface normal prediction, respectively. As shown in Figure 5, DPBridge generates depth maps with sharper object boundaries and more coherent scene geometry, particularly in challenging cases such as urban scenes or complex foreground objects like animals, grass and flowers. In contrast, competing methods like Geowizard and DepthFM often produce over-smoothed outputs or fail to preserve depth discontinuities. In Figure 5, DPBridge demonstrates stronger fidelity in fine-grained geometric details, such as the curvature of architectural structures, the contour of coral and reef, and the fur texture of the cat. Competing methods like Marigold tends to focus on the center object in the image and ignore the environment, while StableNormal produces unreliable outputs. These visual results reinforce the quantitative findings and highlight the effectiveness of DPBridge in modeling structured output distributions with both global coherence and fine details.
Figure 6: Ablation of sampling steps. The performance improves as the number of steps increases, while observing saturation after 10 steps.
@@ 5.3 Ablation Studies
Visual Prior Exploitation. To evaluate the usage of visual priors, we compare finetuning with training from scratch. As shown in the first row of Table 3, finetuning yields significantly better results under identical settings. This is attributed to stronger visual priors from foundation models and faster convergence. Additionally, the last rows of Table 1 and Table 2 show that SD2.1 outperforms SD1.5, highlighting that higher-quality priors further boost performance. These results confirm the advantage of building DPBridge on strong pretrained backbones.
Distribution Alignment Normalization. We next assess the benefit of aligning marginal distributions between the diffusion bridge and standard diffusion processes. As shown in the second and fourth rows of Table 3, adding a distribution alignment normalization step before the noise prediction network consistently improves performance.
Image Consistency Loss. As shown in the third and fourth rows of Table 3, the auxiliary image consistency loss improves performance but has a smaller impact than distribution alignment. This is likely because it mainly affects the boundaries or overlapping regions between masked and valid areas, which are limited in scope. Nonetheless, when combined with other components, it contributes to a notable overall performance gain.
Number of Sampling Steps. We evaluate performance under varying sampling steps during inference. As shown in Figure 6, perception quality improves rapidly with more steps and saturates after 10. Notably, compared to Table 1, where other diffusion-based methods use 50 steps, DPBridge achieves comparable results with significantly fewer steps, highlighting its efficiency.
@@ 6 Conclusion
In this paper, we propose DPBridge, a latent diffusion bridge framework for dense prediction tasks through integrating the strengths of diffusion bridge modeling with visual priors from pretrained foundation models. We derive a fully tractable bridge process to establish direct probability transport in the latent space, and propose finetuning strategies to better adapt pretrained diffusion backbones in bridge construction. While DPBridge achieves competitive performance across different tasks and benchmarks, several limitations remain. First, our method still requires iterative sampling, which makes it slower than feed-forward approaches. Exploring acceleration strategies such as distillation or single-step bridge solvers could further improve practical deployment, especially in real-time settings. Second, our formulation is restricted to certain bridge parameterization, and exploring alternatives to adapt to more advanced pretrained backbone is a promising future direction.
@@ Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4. Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation 5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$. The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Equation 3:
@@
which is the same as in Equation 8. This concludes our derivation.
Appendix B Training and Inference Pipeline
B.1 Pipeline Pseudo-code
Algorithm 1 Training
1:  repeat
2:     Retrieve paired data $(\mathbf{x},\mathbf{y})$
11:  until Converged
Algorithm 2 Inference
1:  $\mathbf{z}_{T}=\mathcal{E}(\mathbf{x})$
2:  for $t=T,\dots,1$ do
7:  end for
8:  $\mathbf{y}=\mathcal{D}(\hat{\mathbf{z}}_{0})$
9:  return $\mathbf{y}$
B.2 Accelerate Inference
Similar to DDIM [45] and DBIM [64], the inference process of DPBridge can be accelerated using non-Markovian sampling scheme that preserves the marginal distributions of intermediate states. Specifically, we perform sampling over a reduced set of timesteps $\{t_{i}\}_{i=1}^{N}$, where $0=t_{0}<t_{1}<\cdots<t_{N-1}<t_{N}=T$, allowing the number of inference steps $N$ to be decoupled from the total number of training steps $T$. The accelerated inference process is given by:
where $g_{n}$ denotes variance parameter that governs the stochasticity of the reverse sampling process. Notably, when $g_{n}=\tilde{\sigma}_{t}^{2}$, as defined Equation 5, the process becomes a Markovian bridge.
Appendix C Network Architecture
The basic structure of our network follows the original design of Stable Diffusion. To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate samples and input image along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by [28], we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
Appendix D Implementation Details
During training, the number of timesteps of is set to be 1000, and we use 50 sampling steps during inference with considerations of both quality and efficiency. Training our method takes 30K iterations using the Adam optimizer with a base learning rate of $3\times 10^{-5}$ and an exponential decay strategy after 100-step warm-up. The batch size is 2, but with 8 steps gradient accumulation, the effective batch size is equivalent to 16. We also apply random horizontal flipping augmentation during training. The whole process takes approximately 3 days to complete on a single Nvidia RTX 3090Ti GPU.
Appendix E Experiments
E.1 Additional Quantitative Results